OpenStudy (anonymous):

limx→0+ (sinx)/(sqrtx) what do?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} sinx / \sqrt{x} = \lim_{x \rightarrow 0} \sqrt{x} (sinx / x) = \lim_{x \rightarrow 0} \sqrt{x} = 0\]

OpenStudy (anonymous):

Since, lim x-> 0 sinx / x = 1

OpenStudy (anonymous):

Thanks man.